Could you please help me find the slope and Y intercept of the line
3x-4y=-24
Thank you for your time and help

Answer by unlockmath(1688)

Hello,
Let's get the equation (3x-4y=-24) into this form: y=mx+b (Note: the b in this equation tells us where the line crosses the y axis and m tells us the slope.)
3x-4y=-24
Subtract 3x and divide by -4 to get:
y=3/4x+6
So now we know the slope is 3/4 and crosses the y axis at 6.
Make sense?
RJ
www.math-unlock.com